OpenStudy (anonymous):

Direct Variation Suppose y varies directly with x. Write a direct variation equation that relates x and y. Then find the value of y when x = 7. 1. y = 8 when x = -4 I'm not asking for the answer, I'd like if someone would explain how to do this because I don't understand it.

OpenStudy (anonymous):

Direct variation means x times some constant is y. So, a direct variation would be of the form: y = kx They give us one point (x,y) = (-4,8). So, y=8 when x=-4. Plug them in: 8 = k(-4) k = -2 Ah, so we finally have our equation: y = kx y = (-2)x And lastly, plug in x=7: y = (-2)(7) = -14

OpenStudy (anonymous):

Thanks so much, thats the answer I got but could someone tell me if I got these two right also? x=7 y=15 when x=6 My answer was "the equation y=2 relates x and y. When x=6, y=2(7), or 14 and x=7 y=3 when x=9 I actually don't have an answer for this one because I got stuck at the dividing part:/

OpenStudy (aripotta):

y = kx substitute: 15 = 6k k = 2.5 y = 7(2.5)

OpenStudy (aripotta):

so when x = 7, y = 17.5

OpenStudy (anonymous):

Oh ok thanks I fixed it. Could you help with the other one?

OpenStudy (aripotta):

y = kx 3 = 9k k = ?

OpenStudy (anonymous):

I'm really confused because my textbook suggest that I divide 3 by 9 but that'd be 0.3. So should I just use 0.3 or flip it and divide 9 by 3?

OpenStudy (aripotta):

3/9 = 1/3 so k = 1/3

OpenStudy (anonymous):

Use 1/3 or 0.3?

OpenStudy (aripotta):

use 1/3

OpenStudy (aripotta):

y = 7(1/3)

OpenStudy (anonymous):

2 1/3?

OpenStudy (aripotta):

yea

OpenStudy (anonymous):

Ok my final answer is "y = 1/3 relates x and y. 1/3(7), or 2 1/3."

OpenStudy (anonymous):

I completely understand now thank you so much @AriPotta

OpenStudy (aripotta):

shouldn't your final answer just be 2 1/3?

OpenStudy (anonymous):

Yeah, but they want me to write a whole sentence

OpenStudy (aripotta):

weird.

OpenStudy (anonymous):

Yup